OpenStudy (accidentalaichan):

I need help with this homework question. Solve the equation 7/8 p - 4 = 10 for p Can someone show steps as well so I can understand it better?

hero (hero):

\[\frac{7}{8}p - 4 = 10\] right?

OpenStudy (accidentalaichan):

@Hero, yes

hero (hero):

Add 4 to both sides..what do you get?

OpenStudy (accidentalaichan):

4 7/8p = 14?

hero (hero):

Where did the extra 4 come from? After adding 4 to sides, you should have (7/8)p = 14 not 4 7/8 p = 14

OpenStudy (accidentalaichan):

Sorry, I'm not good with this question. Thanks.

hero (hero):

So to continue... Multiply the p into the fraction to get \[\frac{7p}{8} = 14\]

hero (hero):

Next... Multiply both sides by 8 7p = (14)(8)

hero (hero):

Don't multiply the 14 and 8 just yet...

hero (hero):

Instead, divide both sides by 7 to get \[p = \frac{(14)(8)}{7}\]

hero (hero):

Then notice that 14/7 = 2

hero (hero):

So re-write as \[p = (2)(8)\]

OpenStudy (accidentalaichan):

So p =16

hero (hero):

Yes...did you understand the steps?

OpenStudy (accidentalaichan):

Yes, thank you so much.

hero (hero):

So that means if I gave you another one, you could solve it in the same manner?